If (an) converges to a and (bn) converges to b show that (an+or - bn) converges to (a+or - b)


1
Expert's answer
2021-03-31T13:54:58-0400

For given ε > 0, there exists a positive integer "N_1" such that

"n> N_1" implies |"a_n-a" | < ε/2.

Moreover, there exists a positive integer "N_2" such that

"n>N_2" implies |"b_n-b" | < ε/2.

Let N := max{"N_1,N_2" }. If n > N, then by the triangle inequality we have

"|(a_n ± b_n) − (a ± b)| ≤ |a_n − a| + |b_n − b| < \dfrac{ε} {2} +\dfrac {ε}{ 2} = ε"

This completes the proof. If "(a_n)_{n=1,2,...}" is a convergent sequence, then the above theorem tells us that

"lim_{n\u2192\u221e} (a_{n+1} \u2212 a_n) = lim_{n\u2192\u221e} a_{n+1} \u2212 lim_{n\u2192\u221e} a_n = 0."